An electron of mass 9.1 * 10^-31 kg is moving in the xy-plane; its speed is 2.0 * 10^5 m/s and its direction of motion makes an angle of 25 degrees with the x axis. What are the components of the momentum of the electron?
Answer #1

momnetum in the x direction = m*vx

=mv cos(25)

=9.1*10^-31*2*10^-5*cos(25)

=9.1*10^-31*2*10^-5*0.906

=1.65*10^-35 Ns

momenutm in the y direciton = mvy

=mvsin(25)

=9.1*10^-31*2*10^-5*0.422

=7.68*10^-36 Ns
Answer #2

mv cos (25 ) along x = 16.49 * 10^-26 along x

mv sin(25) along y = 7.691 *10^-26